On January 1, 2021, the Allegheny Corporation purchased equipment for $199,000. The estimated service life of the equipment is 10 years and the estimated residual value is $12,000. The equipment is expected to produce 248,000 units during its life.

Required:
Calculate depreciation for 2021 and 2022 using each of the following methods.

3. Assume instead the equipment was purchased on October 1, 2021. Calculate depreciation for 2021 and 2022 using each of the two methods. Partial-year depreciation is calculated based on the number of months the asset is in service. (Do not round intermediate answers and round your answers to the nearest whole dollar amount.)

2021 2022
Sum-of-the-years’ digits
One hundred fifty percent declining balance
Answer #1

1.

Sum of the year

Year Cost of the assets Salvage value Depreciable value sum of the year Depreciation
2021 $                 199,000 $           12,000 $                   187,000 1+2+3+4+5+6+7+8+9+10 = 55 $187,000*10/55 = $34,000
2022 $                 199,000 $           12,000 $                   187,000 1+2+3+4+5+6+7+8+9+10 = 55 $187,000*9/55 = $30,600

One hundred fifty percent declining balance

Year - a Net Book value, beginning of year - b Double declained depreciation - c = b/Life of assets*150% Net book value, End of the year - d = b-c
2021 $                 199,000 $                    29,850 $                   169,150
2022 $                 169,150 $                    25,373 $                   143,778

2. Equipment purchased on October 2021

Year Cost of the assets Salvage value Depreciable value sum of the year 2021 Depreciation 2022 Depreciation
2021 $                 199,000 $           12,000 $                   187,000 1+2+3+4+5+6+7+8+9+10 = 55 $187,000*(10/55) * (3 /12) = $8,500 $187,000*(10/55) * (9 /12) = $25,500
2022 $                 199,000 $           12,000 $                   187,000 1+2+3+4+5+6+7+8+9+10 = 55 $187,000*(9/55) * (3 /12) = $7,650
Total depreciation 8500 33150
Depreciation under DDB Method
Year - a Net Book value, beginning of year - b Double declained depreciation - c = b/Life of assets*150% Net book value, End of the year - d = b-c
2021 $                 199,000 $199,000/10*150%*3/12 = $7,463 $                   191,537
2022 $                 191,537 $                                                      28,731 $                   162,806

You can reach me over comment box if you have any doubts. Please rate this answer